OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table? x~ -2 , 0 , 4 f(x)~ 1 , 5 , -59

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=a(x-h)^2+k }\) Plug in each of the given points into this parabola.

OpenStudy (solomonzelman):

For example, \(\color{#000000 }{ \displaystyle \color{blue}{1}=a(\color{red}{-2}-h)^2+k }\)

OpenStudy (anonymous):

I've been doing math all day and I'm literally on the verge of having a breakdown. I'm so tired of this.

OpenStudy (anonymous):

I'm seconds into crying right now. Can you please just work the problem out with me. I know the equation, please just show me how to do it.

OpenStudy (anonymous):

I don't understand why I don't understand math, I'm smart, I know I am but this is just giving me the worst anxiety and I would appreciate your help.

OpenStudy (anonymous):

Ok, so, 1= (-2-h) ^2 +k What do I do next

OpenStudy (solomonzelman):

plug in the other two points into \(\color{#000000 }{ \displaystyle \color{blue}{y}=a(\color{red}{x}-h)^2+k }\)

OpenStudy (anonymous):

so i have two equations?

OpenStudy (solomonzelman):

plug the other points, and show me what equations you get from that...

OpenStudy (anonymous):

I dont understand where im plugging these in. I cant do this im about to freak out im sorry i cant

OpenStudy (solomonzelman):

you are plugging the points into \(\color{#000000 }{ \displaystyle y=a(x-h)^2+k }\)